Prove that W= mgR^2/(R+h)^2


1
Expert's answer
2021-05-04T12:08:30-0400

"W=GMm/R^2=GMm/(R+h)^2=(GM/R^2)\cdot(R^2/(R+h)^2)\cdot m="


"=g\cdot(R^2/(R+h)^2)\cdot m"